OpenStudy (anonymous):

You are given that f(x) = x^2 + 4 and g(x) = 3x - 1 . a.What is the domain & range of f(x) and g(x) b.Find f(x) + g(x) c. What is the domain to your answer to part (b)

OpenStudy (anonymous):

two polynomials, unless otherwise stated domain will be all real numbers

OpenStudy (anonymous):

\[f(x)+g(x)=x^2 + 4 +3x - 1=x^2+3x+3\] domain is still all real numbers

OpenStudy (anonymous):

thanks!